OpenStudy (arbyj):

3x+32 13x+27 how do I find x?

OpenStudy (anonymous):

Are there any equal signs in this?

OpenStudy (arbyj):

i didn't see one, this was on an octagon

OpenStudy (texaschic101):

are these interior or exterior angles of an octagon ?

OpenStudy (arbyj):

exterior

OpenStudy (texaschic101):

Is this a regular octagon or an irregular octagon ?

OpenStudy (arbyj):

regular

OpenStudy (arbyj):



OpenStudy (texaschic101):

I am not the best at geometry, but I do know that each exterior angle of a regular octagon is equal to 45 degrees and when added together equal 360. Thats about all that I know. Sorry...not much help

OpenStudy (arbyj):

thanks anyway (:

OpenStudy (texaschic101):

and a regular octagon has equal sides

OpenStudy (acxbox22):



OpenStudy (texaschic101):

are you looking for perimeter or area ?

OpenStudy (arbyj):

wait i made a mistake its a nonagon

jimthompson5910 (jim_thompson5910):

arbyj, please post a screenshot of the full problem

OpenStudy (arbyj):

i dont if yall can read that but here

OpenStudy (arbyj):

*know

jimthompson5910 (jim_thompson5910):

so it looks like a regular nonagon (9 sided polygon) all 9 sides are congruent if this is true 3x+32 = 13x+27 x = ??

OpenStudy (triciaal):

what did you get for x?

OpenStudy (arbyj):

i got a decimal

OpenStudy (triciaal):

so what decimal did you get for x?

OpenStudy (triciaal):

remember a decimal is just a fraction expressed in units of 10

OpenStudy (arbyj):

.5

OpenStudy (triciaal):

yes 0.5 = 1/2

jimthompson5910 (jim_thompson5910):

0.5 = 5/10 = 1/2 0.5 = five tenths (since the 5 is in the tenths decimal place) I went from 5/10 to 1/2 by dividing both top and bottom by the GCF 5. It's always a good idea to reduce fractions as much as possible

OpenStudy (arbyj):

oh okay

OpenStudy (triciaal):

without a calculator, 3x+32 = 13x+27 -3x -27 from each side 5 = 10 x divide by 10 and reduce x = 5/10 = 1/2